Circular No. 7296

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 U3 (LINEAR) Preliminary parabolic orbital elements by the undersigned from the

observations on IAUC 7295:

T = 1999 Sept.12.27 TT  ω = 58.^{o}82
Ω = 324.90 }2000.0
q = 1.8098 AU  i = 20.58
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 29 3^{h}27.^{m}34 +58^{o}10.^{′}6 1.085 1.905 132.^{o}8 22.^{o}5 16.5
Nov. 3 3 26.69 +58 32.3 1.091 1.926 135.1 21.3 16.5
8 3 25.42 +58 40.2 1.101 1.948 137.2 20.2 16.6
13 3 23.81 +58 34.3 1.114 1.972 139.1 19.2 16.7
18 3 22.13 +58 15.5 1.132 1.998 140.8 18.2 16.8

SUPERNOVA 1999em IN NGC 1637 S. Jha, P. Challis, P. Garnavich, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum was obtained on Oct. 30.34 UT by M. Calkins with the F. L. Whipple Observatory (FLWO) 1.5m telescope (+ FAST spectrograph), revealing SN 1999em to be a type-II event at an early epoch. The spectrum exhibits a very blue continuum with P-Cyg profiles of Hα, Hβ, Hγ, and He I (rest 587.6 nm). The supernova photospheric expansion velocity, measured using the Hβ absorption trough, is 10 300 km/s, adopting the NASA/IPAC Extragalactic Databse recession velocity (717 km/s) for NGC 1637. The spectrum also shows narrow interstellar Na I D (EW = 0.17 nm) at the redshift of NGC 1637, implying that the supernova may be partly obscured by dust. Concurrent images of SN 1999em, taken by K. Stanek with the FLWO 1.2-m telescope, yield a precise position for the supernova of α = 4^{h}41^{m}27^{s}.05, δ = −2^{o}51′45′′.8 (equinox 2000.0, uncertainty ± 0′′.3), which is 15′′.4 west and 17′′.0 south (not north as reported on IAUC 7294) of the host galaxy nucleus.

J. S. Deng, H. L. Cao, D. W. Xu, Y. Lu, H. B. Li, Y. L. Qiu, and J. Y. Hu, Beijing Astronomical Observatory (BAO), report that a low-dispersion spectrum (range 380–740 nm) of SN 1999em was obtained on Oct. 29.7 UT with the BAO 2-m telescope, showing the supernova to be a young type-II event. The spectrum exhibits a very blue continnum with broad emission features of Hα, Hβ, Hγ, Hδ, and Na I D. The narrow Na I D absorption line indicates a recession velocity for NGC 1637 of 710 km/s.

1999 October 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green